# GenOM: Ontology Matching with Description Generation and Large Language Model

### v2

@@ 4.1 Datasets and Evaluation Metrics
The experiments were conducted on the Bio-ML [^6] https://krr-oxford.github.io/OAEI-Bio-ML/2024/index.html track [37] of the Ontology Alignment Evaluation Initiative (OAEI) in 2024, which provides benchmarks specifically designed for biomedical ontology alignment, with a particular focus on machine learning-based systems. This track comprises five sub-tasks involving six widely used biomedical ontologies: Systematized Nomenclature of Medicine Clinical Terms (SNOMED-CT), National Cancer Institute Thesaurus (NCIT), Foundational Model of Anatomy (FMA), Human Disease Ontology (DOID), Orphanet Rare Disease Ontology (ORDO), and Online Mendelian Inheritance in Man (OMIM).
@@ 4.2 Experiment Setup
To systematically evaluate the proposed framework, we designed experimental settings tailored to each module of the pipeline.
@@
Post-processing and Result Fusion: In the final stage, we applied two decision thresholds: the cosine similarity threshold ($\lambda_{\text{cs}}$) and the LLM-based probability threshold ($\lambda_{\text{prob}}$). To determine suitable values, we considered a predefined set of candidate thresholds $\{0.5,0.6,0.7,0.8,0.9\}$ with a step size of $0.1$. This range was chosen to focus on practically meaningful operating points and to exclude extreme high threshold values, which tend to produce overly conservative decisions and poor cross-task robustness in ontology matching. Threshold selection is performed on a single reference task, SNOMED-NCIT (Neoplas). Based on its development results, a unified threshold value of $0.9$ was selected for both $\lambda_{\text{cs}}$ and $\lambda_{\text{prob}}$. These values were then fixed and applied unchanged to all remaining ontology matching tasks. Under such a setting, GenOM may not achieve the optimum performance on other tasks beyond SNOMED-NCIT (Neoplas), but we believe this setting is more consistent with real-world deployment and can reflect the framework’s real performance especially on its robustness on different tasks. In addition, results based on optimal thresholds obtained through task by task tuning are reported in Appendix A, for post-hoc analysis as a reference of GenOM’s optimum performance and for comparison with other systems’ results produced under this optimal setting.
@@
For ease of reference, we use the following abbreviations: Qwen7B (Qwen2.5-7B-Instruct), Qwen14B (Qwen2.5-14B-Instruct), Qwen32B (Qwen2.5-32B-Instruct), and Llama8B (Llama3.1-8B-Instruct). Additionally, we write GenOM (Qwen32B) for the GenOM framework using Qwen32B as the LLM-based alignment model, i.e., the component responsible for predicting equivalence correspondences.
@@ 4.3 Overall Results
Table 5: Overall performance on the five OAEI 2024 Bio-ML tasks. NA indicates no results as the systems do not support the calculation of the metrics. Bold indicates the best performance, while underling indicates the second-best performance. All the baseline results are from the track’s website, where the LogMapLt and the BERTMap series were reimplemented with system submissions, while the other baselines submitted the final result files without system reimplementation, probably using settings optimised for each task. Results of LLM4OM, LogMap-LLM, and Hu et al. were taken from their original papers. LLM4OM and Hu et al. (2025) used fixed thresholds across tasks, while LogMap-LLM relied on OpenAI and Gemini models without an explicit threshold parameter.
Table 5 presented the performance of our method on the Bio-ML track in comparison with several state-of-the-art ontology alignment systems. These systems can broadly be grouped into three categories: (1) traditional ontology matching methods, LogMap [8], LogMapBio [38], LogMapLt [38], and Matcha; (2) pre-trained language model (PLM)-based approaches, BERTMap [15], BERTMapLt, and BioSTransMatch [23]; and (3) LLM-based methods, Hu et al. (2025) [27] (LLaMA-3.3-70B-Instruct), LLM4OM [10] (ChatGPT-3.5), and LogMap-LLM [28] (GPT-4o).
Overall, GenOM (Qwen32B) exhibited competitive performance relative to other baseline systems on the Bio-ML track.
@@
  • On the OMIM–ORDO task, GenOM (Qwen32B) achieved the best performance across all three evaluation metrics. In particular, it outperformed the second-best system on ranking-based metrics, exceeding BERTMap by 0.05 in MRR and by 0.074 in H@1.
@@
Compared with existing LLM-based approaches, GenOM (Qwen32B) achieved strong overall performance while relying on a model with a smaller parameter size. On the two Bio-ML tasks where Hu et al. (2025) [27] reported results (SNOMED–FMA and OMIM–ORDO), GenOM (Qwen32B) increased the F1-score by 0.518 and 0.107, respectively (averaging a 0.313 absolute gain across the two tasks). Across all five Bio-ML tasks, GenOM (Qwen32B) improved the mean F1-score over LLM4OM by 0.205 (35.3%) and over LogMap-LLM by 0.071 (9.88%).
In general, GenOM (Qwen32B) exhibited consistently strong and stable performance in different tasks and evaluation metrics. Even in tasks where it did not achieve the top rank, the performance gaps of the best-performing systems remained small, indicating that GenOM (Qwen32B) operated at a level comparable to state-of-the-art approaches. Furthermore, its strong performance on local ranking metrics demonstrated that GenOM (Qwen32B) was effective in distinguishing semantically similar biomedical concepts, highlighting its effectiveness in fine-grained semantic discrimination.
@@ 5 Definition Generation Analysis
Accurate and informative definitions play an essential role in ontology matching, particularly in the biomedical domain where many concepts differ only subtly [26]. The generated definitions in GenOM are expected to be factually accurate and medically informative, while being sufficiently discriminative to differentiate the target concept from closely related alternatives.
@@ 5.1 Hybrid Evaluation Method
Concretely, for each concept $A$ and its set of generated definitions $\{def_{1},\dots,def_{n}\}$, we first identify a lexically nearest neighbour $B$ from the same ontology using SapBERT [39] embeddings. This neighbour serves as a contrastive reference for evaluating the discriminative quality of definitions. The retrieval relies solely on lexical information (e.g., concept labels), and the nearest neighbour is selected as the top-1 most similar concept within the same ontology. Importantly, generated definitions are not used during this retrieval step, ensuring that the selection of $B$ is independent of generated definitions.
@@ 7.1 Impact of Definition Enrichment on Local Ranking and Candidate Retrieval
We investigated the impact of definition-based semantic enrichment on two stages of the pipeline: LLM-based equivalence judgement and embedding-based candidate generation. For LLM-based equivalence judgement, we compared a label-only setting with a label-plus-definition setting. The local ranking result was shown in Figure 4. In particular, both MRR and Hit@1 showed clear improvements in all tasks except for SNOMED-FMA (body), where performance remains comparable.
@@
For the embedding-based candidate generation stage, we additionally reported Hit@5 and Hit@10, as these metrics are essential for determining the appropriate top-$k$ value—that is, how many candidate concepts should be passed to the LLM for equivalence judgement. As shown in Figure 5, incorporating definition information led to improvements across all three metrics (Hit@1, Hit@5, Hit@10) in all tasks, except for a slight drop in Hit@10 on the SNOMED-NCIT (pharm) task. These results suggested that enriched definitions help retrieve a greater number of correct candidates, thereby increasing the likelihood of including the true match within the top-$k$ shortlist.
@@ 7.2 Effectiveness Compared to Original Exact Matching Methods
As shown in Table 8, GenOM (Qwen32B) surpassed the strongest exact-matching systems on most tasks, with the sole exception of the SNOMED–FMA (Body) alignment. In our earlier workshop version of this work [13], which focused on 7B-level LLMs, we were able to outperform both BERTMapLt and LogMapLt across all tasks by imposing a relatively high cosine-similarity threshold. However, as discussed previously, larger LLMs exhibited much lower sensitivity to the cosine-similarity threshold, with Qwen-32B showing only minor fluctuations in F1 scores between thresholds of 0.6 and 0.9. Under such circumstances, enforcing extremely strict thresholds (e.g. 0.97), while effective for smaller models, is less suitable for larger ones, making a more balanced configuration preferable.
@@ 8 Conclusion, Discussion and Future Work
Appendix A

### v3

@@ 4.1 Datasets and Evaluation Metrics
The experiments were conducted on the Bio-ML [^6] https://krr-oxford.github.io/OAEI-Bio-ML/2024/index.html track [he2022machine] of the Ontology Alignment Evaluation Initiative (OAEI) in 2024, which provides benchmarks specifically designed for biomedical ontology alignment, with a particular focus on machine learning-based systems. This track comprises five sub-tasks involving six widely used biomedical ontologies: Systematized Nomenclature of Medicine Clinical Terms (SNOMED-CT), National Cancer Institute Thesaurus (NCIT), Foundational Model of Anatomy (FMA), Human Disease Ontology (DOID), Orphanet Rare Disease Ontology (ORDO), and Online Mendelian Inheritance in Man (OMIM). These tasks cover multiple biomedical domains, including diseases, anatomical structures, pharmaceuticals, and neoplasms. As summarized in Table 5, the five tasks correspond to different ontology pairs with varying semantic scopes and structural characteristics. All involved ontologies are large-scale biomedical ontologies containing thousands to tens of thousands of concepts, which makes the alignment problem particularly challenging. Evaluating ontology matching systems on such large and heterogeneous ontologies enables a more comprehensive assessment of the scalability and robustness of the proposed approach.
Source Task Category #SrcCls #TgtCls #Ref (equiv)
Mondo OMIM–ORDO Disease 9,648 9,275 3,721
Mondo NCIT–DOID Disease 15,762 8,465 4,686
UMLS SNOMED–FMA Body 34,418 88,955 7,256
UMLS SNOMED–NCIT Pharm 29,500 22,136 5,803
UMLS SNOMED–NCIT Neoplas 22,971 20,247 3,804
Table 5: Statistics of the biomedical ontology matching tasks used in the experiments. #SrcCls and #TgtCls denote the number of source and target concepts, respectively, and #Ref (equiv) indicates the number of reference equivalence mappings.
@@ 4.2 Experiment Setup
The implementation settings are organised according to the pipeline modules of GenOM, with each component corresponding to the stages described in Section 3.
@@
Post-processing and Result Fusion: In the final stage, we applied two decision thresholds: the cosine similarity threshold ($\lambda_{\text{cs}}$) and the LLM-based probability threshold ($\lambda_{\text{prob}}$). To determine suitable values, we considered a predefined set of candidate thresholds $\{0.5,0.6,0.7,0.8,0.9\}$ with a step size of $0.1$. This range was chosen to focus on practically meaningful operating points and to exclude extreme high threshold values, which tend to produce overly conservative decisions and poor cross-task robustness in ontology matching. Threshold selection is performed on a single reference task, SNOMED-NCIT (Neoplas). Based on its development results, a unified threshold value of $0.9$ was selected for both $\lambda_{\text{cs}}$ and $\lambda_{\text{prob}}$. These values were then fixed and applied unchanged to all remaining ontology matching tasks. Under such a setting, GenOM may not achieve the optimum performance on other tasks beyond SNOMED-NCIT (Neoplas), but we believe this setting is more consistent with real-world deployment and can reflect the framework’s real performance especially on its robustness on different tasks. In addition, results based on optimal thresholds obtained through task by task tuning are reported in Appendix A.1, for post-hoc analysis as a reference of GenOM’s optimum performance and for comparison with other systems’ results produced under this optimal setting.
@@
We write GenOM (Qwen32B) for the GenOM framework using Qwen2.5-32B-Instruct as the LLM-based alignment model, i.e. the component responsible for predicting equivalence correspondences.
4.3 Baseline Methods
We compare GenOM with several representative ontology matching systems covering different methodological categories. These include traditional ontology matching systems, LogMap [jimenez2011logmap], LogMapBio [jimenez2015logmap], LogMapLt [jimenez2015logmap], and Matcha; pre-trained language model (PLM)-based approaches including BERTMap [he2022bertmap], BERTMapLt, and BioSTransMatch [menad2023biostransformers]; and recent LLM-based methods, Hu et al. [hu2025matching] (LLaMA-3.3-70B-Instruct), LLM4OM [giglou2024llms4om] (ChatGPT-3.5), and LogMap-LLM [lushnei2025large] (GPT-4o).
Most baseline results are obtained from the official Bio-ML track website. In particular, LogMapLt and the BERTMap series were reimplemented within the track submission framework, while several other systems provided final alignment results without system reimplementation, likely using task-specific optimized configurations. The reported results for LLM4OM, LogMap-LLM, and Hu et al. were taken from their original publications. Among these systems, LLM4OM and Hu et al. use fixed thresholds across tasks, whereas LogMap-LLM relied on OpenAI and Gemini models without an explicit threshold parameter.
@@ 4.4 Overall Results
Table 6: Overall performance on the five OAEI 2024 Bio-ML tasks. NA indicates no results as the systems do not support the calculation of the metrics. Bold indicates the best performance, while underling indicates the second-best performance.
Table 6 presented the performance of our method on the Bio-ML track in comparison with several state-of-the-art ontology alignment systems. Overall, GenOM (Qwen32B) exhibited competitive performance relative to other baseline systems on the Bio-ML track.
@@
  • On the OMIM–ORDO task, GenOM (Qwen32B) achieved the best performance across all three evaluation metrics. In particular, it outperformed the second-best system on ranking-based metrics, exceeding BERTMap by 0.05 in MRR and by 0.074 in H@1. This demonstrates a notable improvement in both metrics.
@@
Compared with existing LLM-based approaches, GenOM (Qwen32B) achieved strong overall performance while relying on a model with a smaller parameter size. On the two Bio-ML tasks where Hu et al. (2025) [hu2025matching] reported results (SNOMED–FMA and OMIM–ORDO), GenOM (Qwen32B) increased the F1-score by 0.518 and 0.107, respectively (averaging a 0.313 absolute gain across the two tasks). Across all five Bio-ML tasks, GenOM (Qwen32B) improved the mean F1-score over LLM4OM by 0.205 (35.3%) and over LogMap-LLM by 0.071 (9.88%).
In general, GenOM (Qwen32B) exhibited consistently strong and stable performance in different tasks and evaluation metrics. Even in tasks where it did not achieve the top rank, the performance gaps of the best-performing systems remained small, indicating that GenOM (Qwen32B) operated at a level comparable to state-of-the-art approaches. Furthermore, its strong performance on local ranking metrics demonstrated that GenOM (Qwen32B) was effective in distinguishing semantically similar biomedical concepts, highlighting its effectiveness in fine-grained semantic discrimination. Representative error cases of the LLM outputs are analysed in Appendix  C.
@@ 5 Definition Generation Analysis
Accurate and informative definitions play an essential role in ontology matching, particularly in the biomedical domain where many concepts differ only subtly [10.14778/3712221.3712222]. The generated definitions in GenOM are expected to be factually accurate and medically informative, while being sufficiently discriminative to differentiate the target concept from closely related alternatives.
@@
For ease of reference, we use the following abbreviations: Qwen7B (Qwen2.5-7B-Instruct), Qwen14B (Qwen2.5-14B-Instruct), Qwen32B (Qwen2.5-32B-Instruct), and Llama8B (Llama3.1-8B-Instruct).
@@ 5.1 Hybrid Evaluation Method
Concretely, for each concept $A$ and its set of generated definitions $\{def_{1},\dots,def_{n}\}$, we first identify a lexically nearest neighbour $B$ from the same ontology using SapBERT [liu2021self] embeddings. This neighbour serves as a contrastive reference for evaluating the discriminative quality of definitions. The retrieval relies solely on lexical information (e.g., concept labels), and the nearest neighbour is selected as the top-1 most similar concept within the same ontology. Importantly, generated definitions are not used during this retrieval step, ensuring that the selection of $B$ is independent of generated definitions.
@@ 7.1 Impact of Definition Enrichment on Local Ranking and Candidate Retrieval
We investigated the impact of definition-based semantic enrichment on two stages of the pipeline: LLM-based equivalence judgement and embedding-based candidate generation.
For LLM-based equivalence judgement, we compared a label-only setting with a label-plus-definition setting. The local ranking result was shown in Figure 4. In particular, both MRR and Hit@1 showed clear improvements in all tasks except for SNOMED-FMA (body), where performance remains comparable. Among these, the improvement is most pronounced in the OMIM–ORDO task, where the performance gains reach 0.18 and 0.235 in terms of MRR and Hit@1, respectively.
For the embedding-based candidate generation stage, we additionally reported Hit@5 and Hit@10, as these metrics are essential for determining the appropriate top-$k$ value—that is, how many candidate concepts should be passed to the LLM for equivalence judgement. As shown in Figure 5, incorporating definition information led to improvements across all three metrics (Hit@1, Hit@5, Hit@10) in all tasks, except for a slight drop in Hit@10 on the SNOMED-NCIT (pharm) task. It is also noteworthy that the inclusion of definitions leads to the most pronounced improvement in Hit@1, whereas the effect becomes considerably attenuated when the evaluation scope is extended to Hit@10, suggesting that definitions primarily enhance fine-grained semantic discrimination rather than broad candidate retrieval.
Overall, these results indicate that incorporating concept definitions improves performance at both stages of the pipeline. In the LLM-based equivalence judgement stage, the additional definition information helps the model better understand the semantics of concepts, which leads to more reliable alignment decisions. In the embedding-based candidate generation stage, definitions provide richer textual information that improves the quality of the retrieved candidates. Together, these findings show that concept definitions support both semantic comparison in the LLM stage and candidate retrieval in the embedding stage. More details can be found in the Appendix A.2
@@ 7.2 Effectiveness Compared to Original Exact Matching Methods
As shown in Table 9, GenOM (Qwen32B) surpassed the strongest exact-matching systems on most tasks, with the sole exception of the SNOMED–FMA (Body) alignment. In our earlier workshop version of this work [song2025genom], which focused on 7B-level LLMs, we were able to outperform both BERTMapLt and LogMapLt across all tasks by imposing a relatively high cosine-similarity threshold. However, as discussed previously, larger LLMs exhibited much lower sensitivity to the cosine-similarity threshold, with Qwen-32B showing only minor fluctuations in F1 scores between thresholds of 0.6 and 0.9. Under such circumstances, enforcing extremely strict thresholds (e.g. 0.97), while effective for smaller models, is less suitable for larger ones, making a more balanced configuration preferable.
@@ 8 Conclusion, Discussion and Future Work
Beyond ontology alignment, the proposed framework can potentially be extended to related semantic matching tasks, such as entity linking, where textual mentions are associated with ontology concepts. While the current work is specifically designed and evaluated for ontology alignment, the combination of definition generation and LLM-based reasoning could, in principle, be adapted to support mention-level semantic matching. We leave this as an interesting direction for future research.
Declarations
Funding. This research received no specific grant from any funding agency in the public, commercial, or not-for-profit sectors.
Data availability. All datasets used in this study are publicly available from open-source ontology repositories.
Conflict of interest. The authors declare that they have no competing interests.
Ethics approval. This study does not involve human participants, animals, or any private or sensitive data and therefore does not require ethical approval.
Appendix A Additional Experimental Details
A.1 Optimal Performance of GenOM(Qwen32B)
@@
A.2 Detailed Ablation Study Results
Table 11: Detailed Results of Definition-based Local Ranking (Qwen32B)
Task MRR Hit@1
OMIM_ORDO (with) 0.903 0.860
OMIM_ORDO (without) 0.723 0.625
NCIT_DOID (with) 0.949 0.924
NCIT_DOID (without) 0.919 0.882
SNOMED_NCIT_pharm (with) 0.918 0.869
SNOMED_NCIT_pharm (without) 0.806 0.739
SNOMED_NCIT_neoplas (with) 0.928 0.888
SNOMED_NCIT_neoplas (without) 0.891 0.841
SNOMED_FMA_body (with) 0.871 0.819
SNOMED_FMA_body (without) 0.866 0.819
Table 12: Detailed Results of embedding-based candidate retrieval (text-embedding-3-small)
Task Hit@1 Hit@5 Hit@10
OMIM_ORDO (with) 0.737 0.879 0.909
OMIM_ORDO (without) 0.724 0.875 0.907
NCIT_DOID (with) 0.886 0.974 0.985
NCIT_DOID (without) 0.786 0.940 0.962
SNOMED_NCIT_pharm (with) 0.747 0.942 0.966
SNOMED_NCIT_pharm (without) 0.712 0.911 0.946
SNOMED_NCIT_neoplas (with) 0.722 0.896 0.932
SNOMED_NCIT_neoplas (without) 0.718 0.896 0.938
SNOMED_FMA_body (with) 0.695 0.912 0.947
SNOMED_FMA_body (without) 0.642 0.883 0.930
@@ B.2 Prompt Template and Analysis
Appendix C Error Analysis of LLM-based Alignment
In this appendix, we present several representative error cases observed in the LLM-based equivalence judgement stage. The analysis focuses on two common error types: false positives, where the model incorrectly predicts equivalence between semantically related but non-equivalent concepts, and false negatives, where the model fails to identify correct equivalence due to differences in conceptual granularity or ontology modelling choices.
It is also important to note that this OM task allows one-to-many and many-to-one mappings. As a result, the result is not restricted to selecting a single target concept for each source concept. This flexibility increases the likelihood of false positive predictions, since multiple semantically related concepts may be identified as potential matches.
C.1 False Positive Case
Source concept Biperiden hydrochloride
Target concept Biperiden
Observation Both concepts share highly similar semantic descriptions. The definitions emphasise that both substances act as anticholinergic agents used to treat symptoms associated with Parkinson’s disease. Their parent classes are also closely related, including ’Antiparkinsonian agent’ and ’Muscarinic cholinergic antagonist’.
Error explanation The inclusion of detailed definitions further strengthens the perceived semantic similarity between the two concepts. As a result, the LLM incorrectly predicts equivalence. However, ’Biperiden hydrochloride’ refers to a specific salt form of the drug, whereas ’Biperiden’ denotes the base compound. Although pharmacologically related, they are modelled as distinct entities in the ontology and should therefore not be aligned as equivalent.
Implication This example illustrates a potential side effect of definition-based semantic enrichment. Additional semantic information may amplify similarities between closely related biomedical concepts, increasing the likelihood of false positive alignments.
Table 14: Representative false positive case in LLM-based ontology alignment.
C.2 False Negative Case
Source concept Quetiapine-containing product
Target concept QUETIAPINE
Observation The source concept refers to a medicinal product category that contains quetiapine as an active ingredient, whereas the target concept denotes the active pharmaceutical ingredient itself. The definitions and hierarchical information highlight that the two concepts belong to different levels of abstraction within the ontology.
Error explanation From a semantic perspective, the model’s behaviour is understandable. One concept represents a class of medicinal products containing a particular ingredient, while the other represents the chemical substance itself. As a result, the LLM tends to interpret them as related but not strictly equivalent concepts.
Implication According to the ground truth alignment provided in the benchmark dataset, these two concepts are considered equivalent. Therefore, the model prediction is counted as a false negative under the evaluation setting. This example suggests that some evaluation errors may arise from differences in conceptual granularity or ontology modelling choices rather than genuine semantic misunderstandings by the model.
Table 15: Representative false negative case in LLM-based ontology alignment.